Jane travels 6 times faster than Brenda. Traveling in opposite directions, they are 210 miles apart after 2 hours. Find their rates of trave
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52879)

.
Jane travels 6 times faster than Brenda. Traveling in opposite directions, they are 210 miles apart after 2 hours. Find their rates of trave
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Brenda: 15 mph,
Jane  : 90 mph.


Answer by Alan3354(69443)

Jane travels 6 times faster than Brenda. Traveling in opposite directions, they are 210 miles apart after 2 hours. Find their rates of travel
------------------
6 times faster = 7 times as fast.
---
The sum of their speeds is 210/2 = 105 mi/hr
---
J + B = 105
J = 7B
---
7B + B = 105
Brenda's speed = 105/8 mi/hr
Jane's speed = 735/8 mi/hr
======================================
100% faster is 1 time faster = 2x as fast.
200% faster is 2 times faster = 3x as fast.
300% faster is 3 times faster = 4x as fast.